# TORAX: A Fast and Differentiable Tokamak Transport Simulator in JAX

### v2

@@ II.1 Governing equations
    $\frac{16\pi^{2}\sigma_{||}\mu_{0}\hat{\rho}\Phi_{b}^{2}}{F^{2}}\left(\frac{\partial\psi}{\partial t}-\frac{\hat{\rho}\dot{\Phi}_{b}}{2\Phi_{b}}\frac{\partial\psi}{\partial\hat{\rho}}\right)=\\ \frac{\partial}{\partial\hat{\rho}}\left(\frac{g_{2}g_{3}}{\hat{\rho}}\frac{\partial\psi}{\partial\hat{\rho}}\right)-\frac{8\pi^{2}V^{\prime}\mu_{0}\Phi_{b}}{F^{2}}\langle j_{ni}\rangle$ (4)
where $T_{i,e}$ are ion and electron temperatures, $n_{i,e}$ are ion and electron densities, and $\psi$ is poloidal flux. $\chi_{i,e}$ are ion and electron heat conductivities, $q_{i,e}^{\mathrm{conv}}$ are ion and electron heat convections, $D_{e}$ is electron particle diffusivity, and $V_{e}$ is electron particle convection. $Q_{i,e}$ are the total ion and electron heat sources, and $S_{n}$ is the total electron particle source. $V^{\prime}\equiv dV/d\hat{\rho}$, i.e. the flux surface volume derivative. $\sigma_{||}$ is the plasma neoclassical conductivity, and $\langle j_{ni}\rangle$ is the flux-surface-averaged non-inductive current, being comprised of the bootstrap current and external current drive. $F\equiv RB_{\varphi}$ is the toroidal field flux function, where $R$ is major radius and $B_{\varphi}$ the toroidal magnetic field. $\Phi_{b}$ is the toroidal flux enclosed by the last closed flux surface, and $\dot{\Phi}_{b}$ is its time derivative, non-zero with time-varying toroidal magnetic field and/or last closed flux surface shape. $\mu_{0}$ is the permeability of free space. The geometric quantities $g_{0}$, $g_{1}$, $g_{2}$ and $g_{3}$ are defined as follows.
@@
See Ref. [23] for a more comprehensive summary and derivation of the governing equations and associated geometric quantities. Future work can extend the governing equations to include momentum transport, and multiple ion species including impurities. Details of the physics models underlying the PDE coefficients is provided in section III.
@@ II.3.6 Timestep ($\Delta t$) calculation
The adaptive timestep method protects against traversing through fast transients in the simulation, by enforcing $\Delta t\propto\chi$.

### v3

@@ II.1 Governing equations
    $\frac{16\pi^{2}\sigma_{||}\mu_{0}\hat{\rho}\Phi_{b}^{2}}{F^{2}}\left(\frac{\partial\psi}{\partial t}-\frac{\hat{\rho}\dot{\Phi}_{b}}{2\Phi_{b}}\frac{\partial\psi}{\partial\hat{\rho}}\right)=\\ \frac{\partial}{\partial\hat{\rho}}\left(\frac{g_{2}g_{3}}{\hat{\rho}}\frac{\partial\psi}{\partial\hat{\rho}}\right)-\frac{8\pi^{2}V^{\prime}\mu_{0}\Phi_{b}}{F^{2}}\langle\mathbf{B}\cdot\mathbf{j}_{ni}\rangle$ (4)
where $T_{i,e}$ are ion and electron temperatures, $n_{i,e}$ are ion and electron densities, and $\psi$ is poloidal flux. $\chi_{i,e}$ are ion and electron heat conductivities, $q_{i,e}^{\mathrm{conv}}$ are ion and electron heat convections, $D_{e}$ is electron particle diffusivity, and $V_{e}$ is electron particle convection. $Q_{i,e}$ are the total ion and electron heat sources, and $S_{n}$ is the total electron particle source. $V^{\prime}\equiv dV/d\hat{\rho}$, i.e. the flux surface volume derivative. $\sigma_{||}$ is the plasma neoclassical conductivity, and $\langle\mathbf{B}\cdot\mathbf{j}_{ni}\rangle$ is the flux-surface-averaged non-inductive current (comprised of the bootstrap current and external current drive) projected onto and multiplied by the magnetic field. $F\equiv RB_{\varphi}$ is the toroidal field flux function, where $R$ is major radius and $B_{\varphi}$ the toroidal magnetic field. $\Phi_{b}$ is the toroidal flux enclosed by the last closed flux surface, and $\dot{\Phi}_{b}$ is its time derivative, non-zero with time-varying toroidal magnetic field and/or last closed flux surface shape. $\mu_{0}$ is the permeability of free space. The geometric quantities $g_{0}$, $g_{1}$, $g_{2}$ and $g_{3}$ are defined as follows.
@@
See Ref. [23] for a more comprehensive summary and derivation of the governing equations and associated geometric quantities. See Appendix A for a breakdown in how the $\dot{\Phi}_{b}$ term leads to additional convective and source terms. Future work can extend the governing equations to include momentum transport, and multiple ion species including impurities. Details of the physics models underlying the PDE coefficients is provided in section III.
@@ II.3.6 Timestep ($\Delta t$) calculation
The adaptive timestep method protects against traversing through fast transients in the simulation, by enforcing $\Delta t\propto 1/\chi$.
@@ VII Acknowledgements
Appendix A Time dependent geometry impact on convective and source terms
For ion and electron heat transport, suppressing the species subscript and multiplying the equation by $V^{\prime}$, we obtain
$\frac{3}{2}V^{\prime-2/3}\left(\frac{\partial}{\partial t}-\frac{\dot{\Phi}_{b}}{2\Phi_{b}}\frac{\partial}{\partial\hat{\rho}}\hat{\rho}\right)\left[V^{\prime 5/3}nT\right]=\\ \frac{\partial}{\partial\hat{\rho}}\left[\chi n\frac{g_{1}}{V^{\prime}}\frac{\partial T}{\partial\hat{\rho}}-g_{0}q^{\mathrm{conv}}T\right]+V^{\prime}Q$ (42)
Moving the $\dot{\Phi}_{b}$ term to the RHS, we obtain:
$\frac{3}{2}V^{\prime-2/3}\frac{\partial}{\partial t}\left(V^{\prime 5/3}nT\right)=\frac{3}{4}V^{\prime-2/3}\frac{\dot{\Phi}_{b}}{\Phi_{b}}\frac{\partial}{\partial\hat{\rho}}\hat{\rho}V^{\prime 5/3}nT\\ +\frac{\partial}{\partial\hat{\rho}}\left[\chi n\frac{g_{1}}{V^{\prime}}\frac{\partial T}{\partial\hat{\rho}}-g_{0}q^{\mathrm{conv}}T\right]+V^{\prime}Q$ (43)
Utilizing the identity $a\frac{\partial(bT)}{\partial\hat{\rho}}=\frac{\partial(abT)}{\partial\hat{\rho}}-bT\frac{\partial a}{\partial\hat{\rho}}$, and rearranging terms, we obtain:
$\frac{3}{2}V^{\prime-2/3}\frac{\partial}{\partial t}\left(V^{\prime 5/3}nT\right)=\\ \frac{\partial}{\partial\hat{\rho}}\left[\chi n\frac{g_{1}}{V^{\prime}}\frac{\partial T}{\partial\hat{\rho}}-\left(g_{0}q^{\mathrm{conv}}-\frac{3}{4}\frac{\dot{\Phi}_{b}}{\Phi_{b}}\hat{\rho}V^{\prime}n\right)T\right]+\\ V^{\prime}Q+\frac{1}{2}V^{\prime\prime}\frac{\dot{\Phi}_{b}}{\Phi_{b}}\hat{\rho}nT$ (44)
Non-zero $\dot{\Phi}_{b}$ thus results in an effective convection term equal to -$\frac{3}{4}\frac{\dot{\Phi}_{b}}{\Phi_{b}}\hat{\rho}V^{\prime}n$, and an effective additional source term equal to $\frac{1}{2}V^{\prime\prime}\frac{\dot{\Phi}_{b}}{\Phi_{b}}\hat{\rho}nT$, where $V^{\prime\prime}$ is the second derivative of the plasma volume with respect to $\hat{\rho}$.
For particle transport, using similar considerations, we obtain the following governing equation:
$\frac{\partial}{\partial t}\left(n_{e}V^{\prime}\right)=\\ \frac{\partial}{\partial\hat{\rho}}\left[D_{e}n_{e}\frac{g_{1}}{V^{\prime}}\frac{\partial n_{e}}{\partial\hat{\rho}}-\left(g_{0}V_{e}-\frac{\dot{\Phi}_{b}}{2\Phi_{b}}\hat{\rho}V^{\prime}\right)n_{e}\right]+V^{\prime}S_{n}$ (45)
where the $\dot{\Phi}_{b}$ terms leads to an additional effective convection term of -$\frac{\dot{\Phi}_{b}}{2\Phi_{b}}\hat{\rho}V^{\prime}$, and no additional effective source terms.
For current diffusion, we obtain:
$\frac{16\pi^{2}\sigma_{||}\mu_{0}\hat{\rho}\Phi_{b}^{2}}{F^{2}}\frac{\partial\psi}{\partial t}-\frac{\hat{\rho}\dot{\Phi}_{b}}{2\Phi_{b}}\frac{\partial\psi}{\partial\hat{\rho}}=\\ \frac{\partial}{\partial\hat{\rho}}\left(\frac{g_{2}g_{3}}{\hat{\rho}}\frac{\partial\psi}{\partial\hat{\rho}}+8\pi^{2}\mu_{0}\Phi_{b}\dot{\Phi}_{b}\frac{\sigma_{||}\hat{\rho}^{2}}{F^{2}}\psi\right)-\\ \frac{8\pi^{2}V^{\prime}\mu_{0}\Phi_{b}}{F^{2}}\langle\mathbf{B}\cdot\mathbf{j}_{ni}\rangle-8\pi^{2}\mu_{0}\Phi_{b}\dot{\Phi}_{b}\psi\frac{\partial}{\partial\hat{\rho}}\left(\frac{\sigma_{||}\hat{\rho}^{2}}{F^{2}}\right)$ (46)
where the $\dot{\Phi}_{b}$ terms leads to an additional effective convection term of - $8\pi^{2}\mu_{0}\Phi_{b}\dot{\Phi}_{b}\frac{\sigma_{||}\hat{\rho}^{2}}{F^{2}}$, and an additional effective source term of $-8\pi^{2}\mu_{0}\Phi_{b}\dot{\Phi}_{b}\psi\frac{\partial}{\partial\hat{\rho}}\left(\frac{\sigma_{||}\hat{\rho}^{2}}{F^{2}}\right)$.